OpenStudy (anonymous):

Given z = 2 - j2 is the root of 2z^3 – 9z^2 +20z – 8 = 0 find the remaining roots of the equation.

OpenStudy (mathmate):

Since the conjugate is another root, we have z=2+j2, which means that (z+2+j2)(z+2-j2)=(z^2+4z+8) is a factor. The remaining factor can be deduced by synthetic division (2z-1), or z=1/2.